A university’s administrator proposes to do an analysis of the proportion of graduates who have not found employment in their major field one year after graduation. In previous years, the percentage averaged 12%. How many graduates must be sampled to be 92% confident that your sample percentage has a margin of error of two percentage points?

if it can draw graph i need a graph too and i need clear write thank you for help
Answer #1

The sample size formula is given as below:

n = p*q*(Z/E)^2

We are given

p = 0.12

q = 1 – p = 0.88

Confidence level = 92%

Critical Z value = 1.7507

(by using z-table)

Margin of error = E = 0.02

The sample size is given as below:

n = p*q*(Z/E)^2

n = 0.12*0.88*(1.7507/0.02)^2

n = 809.1469

Required sample size = 810

Answer: 810